SsTDOGRednek:

The area of a circle is 20cm^2. What is the circumference in centimeters? Express in terms of Pie.

SsTDOGRednek:

Could anyone help me with this? I don't understand it. I searched it up and I still don't understand it.

Indagurity:

Isn't the equation for circumference like something involving pi

SsTDOGRednek:

Yeah, it would be a number then the PI symbol after it

kristen000021:

Using the formulas A=πr2 C=2πr Solving for C C=2√(πA)=2·√(π·20)≈15.85331cm C=15.85 cm

SsTDOGRednek:

I sure hope this is right because I have to put these down for noted lol

SsTDOGRednek:

I'll double check it with teh calculator and then the teacher just to make sure because I need to have the right notes

jhonyy9:

@kristen000021 wrote:
Using the formulas A=πr2 C=2πr Solving for C C=2√(πA)=2·√(π·20)≈15.85331cm C=15.85 cm
from formula of area you need calcule the r \[A=pir^2 \] \[r=\sqrt{A/\pi }\] and now you can use this value of r to calcule the circumference

SsTDOGRednek:

r is radius right?

kristen000021:

yea

SsTDOGRednek:

Alr, good to know. I need to remember that.